18) A moderate wind accelerates a pebble over a horizontal xy plane with a constant acceleration ModifyingAbove a With right-arrow equals left-parenthesis 5.0ModifyingAbove i With caret plus 7.0ModifyingAbove j With caret right-parenthesis m/s squared . At time t = 0, the velocity is left-parenthesis 5.0ModifyingAbove i With caret right-parenthesis m/s . What are the (a)(significant digits of 3) magnitude and (b)(significant digits of 3) angle of its velocity when it has been displaced by 12.0 m parallel to the x axis?

Please include units
Answer #1

Acceleration of pebble is a = (5.02 + 7.09) m/s2 At time t = 0, velocity of pebble is Ū; = 5.02 m/s ā=azî + ay] =(5.0 +7.09)=